# High Probability Bound for Cross-Learning Contextual Bandits with Unknown Context Distributions

### v1

@@ Abstract
Technically, the most interesting setting for the cross-learning contextual bandits problem is when the losses are chosen adversarially but the contexts are i.i.d. samples from an unknown distribution $\nu$. Recently, Schneider and Zimmert (2023) gave an algorithm achieving nearly optimal $\tilde{O}(\sqrt{KT})$ regret in this scenario.
@@
In this paper, we show that the algorithm indeed achieves nearly optimal $\tilde{O}(\sqrt{KT})$ regret with high-probability by conducting a more in-depth analysis. Our analysis introduces several new insights. In particular, we exploit the weak dependency structure between different epochs, which was overlooked in previous work. One difficulty of doing so is that standard martingale inequalities are not directly applicable, so we refine martingale inequalities to complete our analysis.
@@ 1.1 Related Works
The cross-learning contextual bandits problem was first proposed in Balseiro et al. (2019). They achieve the nearly optimal $\tilde{O}(\sqrt{KT})$ regret under two scenarios: (1) when both losses and contexts are stochastic, and (2) when losses are adversarial and contexts are stochastic with a known distribution. When losses are adversarial and contexts are stochastic with an unknown distribution, they only achieve the suboptimal $\tilde{O}(K^{1/3}T^{2/3})$ regret. More recently, Schneider and Zimmert (2023) gave a new algorithm that achieves the nearly optimal $\tilde{O}(\sqrt{KT})$ regret in expectation under adversarial losses and stochastic contexts with an unknown distribution.
@@
Another interesting application of the cross-learning contextual bandits problem is the sleeping bandits problem (Kleinberg et al., 2010; Neu and Valko, 2014; Kale et al., 2016; Saha et al., 2020). In this problem, a certain set of arms is unavailable in each round. The sleeping bandits problem is motivated by instances like some items might go out of stock in retail stores or on a certain day some websites could be down. When losses are adversarial and availabilities are stochastic, previous work either requires exponential computing time (Kleinberg et al., 2010; Neu and Valko, 2014) or results in suboptimal regret (Kale et al., 2016; Saha et al., 2020). The first computationally efficient algorithm with optimal regret $\tilde{O}(\sqrt{KT})$ is proposed in Schneider and Zimmert (2023) by modeling the problem as a cross-learning contextual bandit.
@@
2 Preliminaries
2.1 Problem Statement
We study a contextual $K$-armed bandit problem over $T$ rounds, with contexts belonging to the set $[C]$. At the beginning of the problem, an oblivious adversary selects a sequence of losses $\ell_{t,c}(k)\in[0,1]$ for every round $t\in[T]$, every context $c\in[C]$, and every arm $k\in[K]$. In each round $t$, we begin by sampling a context $c_{t}\sim\nu$ i.i.d. from an unknown distribution $\nu$ over $[C]$, and we reveal this context to the learner. Based on this context, the learner selects an arm $a_{t}\in[K]$ to play. The adversary then reveals the function $\ell_{t,c}(a_{t})$, and the learner suffers loss $\ell_{t,c_{t}}\left(a_{t}\right)$. Notably, the learner observes the loss for every context $c\in[C]$, but only for the arm $a_{t}$ they actually played.
@@ 2.1 Problem Statement
Schneider and Zimmert (2023) designed an algorithm that achieves an expected regret bound of
$\mathbb{E}\left[\operatorname{\mathrm{Reg}}(\pi)\right]\leq\tilde{O}(\sqrt{KT})$
for any policy $\pi$. We will show that the algorithm in Schneider and Zimmert (2023) actually provides a high-probability regret bound. That is, we will prove the following result.
Theorem 1.
For $\iota=2\log(8KT)$, $L=\sqrt{\frac{\iota KT}{\log(K)}}=\widetilde{\Theta}(\sqrt{KT})$, $\gamma=\frac{16\iota}{L}=\widetilde{\Theta}(1/\sqrt{KT})$, and $\eta=\frac{\gamma}{2(2L\gamma+\iota)}=\widetilde{\Theta}(1/\sqrt{KT})$, the algorithm in Schneider and Zimmert (2023) yields a regret bound of $\operatorname{\mathrm{Reg}}(\pi)=\widetilde{\Theta}(\sqrt{KT})$ with high probability for any policy $\pi$.
2.2 Existing Analysis
@@
$p_{t,c}=\operatorname*{arg\,min}_{p\in\Delta([K])}\left\langle p,\sum_{s=1}^{t-1}\hat{\ell}_{s,c}\right\rangle-\frac{1}{\eta}F(p)$
@@
for each context $c$, where $F(p)=\sum_{i=1}^{K}p_{i}\log(p_{i})$ is the unnormalized negative entropy, $\eta$ is a learning rate, and $\hat{\ell}$ are loss estimates to be defined later. The algorithm then essentially samples the action $a_{t}$ to be played in round $t$ from distribution $p_{t,c_{t}}$.
@@
The key novelty in Schneider and Zimmert (2023) lies in the construction of the loss estimates $\hat{\ell}$. An intuitive construction is defined as follows:
^$\hat{\ell}_{t,c}(a)=\frac{\ell_{t,c}(a)}{\mathbb{E}_{c\sim\nu}[p_{t,c}(a)]}\operatorname{\mathbbm{1}}(a_{t}=a).$
@@
That is, it uses the classic importance-weighted estimator with $\mathbb{E}_{c\sim\nu}[p_{t,c}(a)]$ as the importance. A straightforward analysis shows that this estimator yields a regret bound of $\tilde{O}(\sqrt{KT})$. However, the denominator term $\mathbb{E}_{c\sim\nu}[p_{t,c}(a)]$ is uncomputable because we do not know the distribution of contexts $\nu$. One may attempt to circumvent this issue by replacing the expected importance $\mathbb{E}_{c\sim\nu}[p_{t,c}(a)]$ with the empirical importance $\frac{1}{t}\sum_{s=1}^{t}p_{t,c_{s}}(a)$. It is not hard to see that whether we achieve the desired $\tilde{O}(\sqrt{KT})$ regret depends on how well the empirical importance weight $\frac{1}{\frac{1}{t}\sum_{s=1}^{t}p_{t,c_{s}}(a)}$ concentrates around the expected importance weight $\frac{1}{\mathbb{E}_{c\sim\nu}[p_{t,c}(a)]}$. However, the empirical importance weight $\frac{1}{\frac{1}{t}\sum_{s=1}^{t}p_{t,c_{s}}(a)}$ may not concentrate well around the expected importance weight $\frac{1}{\mathbb{E}_{c\sim\nu}[p_{t,c}(a)]}$. This is because the probability vector $p_{t,c}$ is not independent of the previous contexts $c_{s}$, which makes standard concentration inequalities inapplicable.
@@
To address this difficulty, Schneider and Zimmert (2023) divides the time horizon into epochs of equal length $L$. At the end of each epoch $e$, the algorithm stores the FTRL distribution at the current time $t=eL$ in a new distribution $s_{e}$; that is, it takes $s_{e,c}(a)=p_{t,c}(a)$ for each context $c$ and each arm $a$. The algorithm further decouples the distribution played by the algorithm and the distribution used to estimate the loss vector. For each time $t$ in epoch $e+2$, the algorithm observes the loss $\ell_{t,c}(a)$ for each arm $a$ and context $c$ with probability $f_{e}(a)\triangleq\mathbb{E}_{c\sim\nu}[s_{e,c}(a)/2]$. The algorithm then estimates the expected importance $f_{e}(a)$ using an empirical importance $\hat{f}_{e}(a)$ constructing solely from contexts in epoch $e+1$. Finally, the algorithm constructs ^$\hat{\ell}_{t,c}(a)$ as an importance-weighted estimator with $\hat{f}_{e}(a)$ serving as the importance.
The empirical importance weight $\frac{1}{\hat{f}_{e}(a)}$ concentrates around the expected importance weight $\frac{1}{f_{e}(a)}$ because the algorithm only uses samples from epoch $e+1$ to construct the estimator, and these samples are independent of the estimand $\frac{1}{f_{e}(a)}$. However, since the algorithm uses samples from only a single epoch rather than the entire time horizon, the concentration $|\frac{1}{\hat{f}_{e}(a)}-\frac{1}{f_{e}(a)}|$ is not tight enough. To achieve the desired $\tilde{O}(\sqrt{KT})$ regret under a not tight enough concentration, Schneider and Zimmert (2023) bounds only the expected bias of importance estimator $\mathbb{E}[\frac{1}{\hat{f}_{e}(a)}-\frac{1}{f_{e}(k)}]$ rather than providing a high-probability bias bound. Bounding only the expected bias gives a small enough bound, however, they can achieve a bound only on the expected regret from a bound on the expected bias.
2.3 Our Techniques
We overcome this difficulty and show that their algorithm actually achieves a high-probability bound. Our key observation is that different epochs in their algorithm are only weakly dependent on each other. Thus, the bias $\frac{1}{\hat{f}_{e}(a)}-\frac{1}{f_{e}(a)}$ for each epoch $e$ is also only weakly dependent on each other. Therefore, although we cannot establish a small enough bound for the bias of a single epoch $\frac{1}{f_{ek}}-\frac{1}{\widehat{f}_{ek}}$, we can give a small enough bound for the cumulative bias across all epochs $\sum_{e}\frac{1}{f_{ek}}-\frac{1}{\widehat{f}_{ek}}$. We then use the bound on the cumulative bias to bound the cumulative regret.
In addition to utilizing the weak dependency structure between different epochs, we also address two further technical difficulties to establish our result. The first difficulty is that the existing regret decomposition is too crude to yield a high-probability bound. Schneider and Zimmert (2023) establish an $\tilde{O}(\sqrt{KT})$ expected regret by decomposing the regret into different parts and bounding each part separately. Although their decomposition gives an $\tilde{O}(\sqrt{KT})$ expected regret bound, it is too crude to derive a tight high-probability regret bound, even after utilizing the weak dependency structure. We carefully rearrange the regret decomposition to address this difficulty.
Secondly, we cannot simply apply standard martingale concentration inequalities to $\sum_{e}\frac{1}{f_{ek}}-\frac{1}{\widehat{f}_{ek}}$ to bound its deviation. The main problem is that the random variable $\frac{1}{f_{ek}}-\frac{1}{\widehat{f}_{ek}}$ is not almost surely bounded by a constant, which makes standard martingale concentration inequalities inapplicable. We introduce a surrogate sequence of random variables as a bridge to address this problem. We bound the sum over the surrogate sequence, and show that the sum over the real sequence is equal to the surrogate sequence with high probability.
3 Main Analysis
3.1 The Algorithm in Schneider and Zimmert (2023)
@@
In this subsection, we briefly recall the algorithm proposed in Schneider and Zimmert (2023) for the sake of completeness. Readers familar with Schneider and Zimmert (2023) can skip this subsection safely.
@@
$p_{t,c}=\operatorname*{arg\,min}_{p\in\Delta([K])}\left\langle p,\sum_{s=1}^{t-1}\hat{\ell}_{s,c}\right\rangle-\frac{1}{\eta}F(p)$
for each context $c$, where $F(p)=\sum_{i=1}^{K}p_{i}\log(p_{i})$ is the unnormalized negative entropy, $\eta$ is the learning rate, and $\hat{\ell}$ are loss estimates to be defined later. The algorithm will not sample the action $a_{t}$ played in round t directly from $p_{t}$ but from a distribution $q_{t}$ to be defined later.
To construct loss estimates $\hat{\ell}$, the algorithm divides the time horizon into epochs of equal length $L$. We let $\mathcal{T}_{e}$ to denote the set of rounds in the $e$-th epoch. At the end of each epoch, the algorithm takes a single snapshot of the underlying FTRL distribution $p_{t}$ for each context and arm. That is, the algorithm takes
$s_{e+2,c}(a)=p_{eL,c}(a),\text{ where }s_{1,c}(a)=s_{2,c}(a)=\begin{cases}\frac{1}{\left|\mathcal{A}_{c}\right|}&\text{ if }a\in\mathcal{A}_{c}\\ 0&\text{ otherwise.}\end{cases}$
@@
^$\widehat{\ell}_{t}(a)=\frac{2\ell_{t}(a)}{\widehat{f}_{e}(a)+\frac{3}{2}\gamma}\operatorname{\mathbbm{1}}\left(A_{t}=a\wedge S_{t}\wedge t\in\mathcal{T}_{e}^{\ell}\right)$

### v2

@@ 1 Introduction
Technically, the most interesting setting for the cross-learning contextual bandits problem is when the losses are chosen adversarially but the contexts are i.i.d. samples from an unknown distribution $\nu$. Recently, Schneider and Zimmert (2023) gave an algorithm achieving nearly optimal $\widetilde{O}(\sqrt{KT})$ expected regret in this scenario.
@@
In this paper, we show that the algorithm indeed achieves nearly optimal $\widetilde{O}(\sqrt{KT})$ regret with high-probability. The key contribution of our paper is the following theorem.
Theorem 1 (Informal).
The algorithm in Schneider and Zimmert (2023) yields a regret bound of order $\widetilde{O}(\sqrt{KT})$ with high probability for any policy $\pi$.
In this section we only give the informal version of Theorem 1. The formal version can be found in Section 4.
1.1 Technical Overview
Our theorem is built on a new and more in-depth analysis of the algorithm in Schneider and Zimmert (2023). This new analysis introduces several new insights. In particular, we exploit the weak dependency structure between different epochs, which was overlooked in previous work. One difficulty of doing so is that standard martingale inequalities are not directly applicable, so we refine martingale inequalities to complete our analysis.
To prepare the readers for our new analysis, we first briefly introduce the algorithm in Schneider and Zimmert (2023). The algorithm in Schneider and Zimmert (2023) is an EXP3-type algorithm. The key novelty in their algorithm is the construction of the loss estimates $\widehat{\ell}$ used in the FTRL subroutine. Due to some technical problems we detail later, the algorithm decomposes the time horizon into epochs of equal length. In each epoch $e$, the algorithm first estimates the probability [^1] For technical reasons, in the actual algorithm, the value $f_{e}(a)$ actually represents the probability of observing the reward of each arm $a$ in epoch $e+2$. For ease of understanding, here we instead let it represent the probability of observing the reward of each $\operatorname{arm}a$ in each epoch $e$. $f_{e}(a)$ of observing the reward of each arm $a$ in epoch $e$ by an estimator $\widehat{f}_{e}(a)$, which is constructed exclusively from samples in epoch $e-1$. Note that thanks to the cross-learning structure, the probability of observing the reward of each arm $a$ is independent of the contexts. The algorithm then constructs the loss estimates as an importance-weighted estimator with $\frac{1}{\widehat{f}_{e}(a)}$ as the importance weight.
Schneider and Zimmert (2023) showed that the performance of the algorithm depends on how well the empirical importance weight $\frac{1}{\widehat{f}_{e}(a)}$ concentrates around the expected importance weight $\frac{1}{f_{e}(a)}$. Since the estimator $\widehat{f}_{e}(a)$ is constructed exclusively from samples in a single epoch rather than the entire time horizon, the concentration $|\frac{1}{f_{e}(a)}-\frac{1}{\widehat{f}_{e}(a)}|$ is not tight enough. To achieve the desired $\widetilde{O}(\sqrt{KT})$ regret under a not tight enough concentration, Schneider and Zimmert (2023) bounds only the expected bias of importance estimator $\mathbb{E}[\frac{1}{f_{e}(a)}-\frac{1}{\widehat{f}_{e}(a)}]$ rather than providing a high-probability bias bound. Bounding only the expected bias gives a small enough bound, however, they can achieve a bound only on the expected regret from a bound on the expected bias.
We overcome this difficulty and show that their algorithm actually achieves a high-probability bound. Our key observation is that different epochs in their algorithm are only weakly dependent on each other. Thus, the bias $\frac{1}{\widehat{f}_{e}(a)}-\frac{1}{f_{e}(a)}$ for each epoch $e$ is also only weakly dependent on each other. Therefore, although we cannot establish a small enough bound for the bias of a single epoch $\frac{1}{f_{e}(a)}-\frac{1}{\widehat{f}_{e}(a)}$, we can give a small enough bound for the cumulative bias across all epochs $\sum_{e}\frac{1}{f_{e}(a)}-\frac{1}{\widehat{f}_{e}(a)}$. We then use the bound on the cumulative bias to bound the cumulative regret.
In addition to utilizing the weak dependency structure between different epochs, we also address two further technical difficulties to establish our result. The first difficulty is that the existing regret decomposition is too crude to yield a high-probability bound. Schneider and Zimmert (2023) establish an $\widetilde{O}(\sqrt{KT})$ expected regret by decomposing the regret into different parts and bounding each part separately. Although their decomposition gives an $\widetilde{O}(\sqrt{KT})$ expected regret bound, it is too crude to derive a tight high-probability regret bound, even after utilizing the weak dependency structure. We carefully rearrange the regret decomposition to address this difficulty.
Secondly, we cannot simply apply standard martingale concentration inequalities to $\sum_{e}\frac{1}{f_{e}(a)}-\frac{1}{\widehat{f}_{e}(a)}$ to bound its deviation. The main problem is that the random variable $\frac{1}{f_{e}(a)}-\frac{1}{\widehat{f}_{e}(a)}$ is not almost surely bounded by a constant, which makes standard martingale concentration inequalities inapplicable. We introduce a surrogate sequence of random variables as a bridge to address this problem. We bound the sum over the surrogate sequence, and show that the sum over the real sequence is equal to the surrogate sequence with high probability.
@@ 1.2 Related Works
The cross-learning contextual bandits problem was first proposed in Balseiro et al. (2019). They achieve the nearly optimal $\widetilde{O}(\sqrt{KT})$ regret under two scenarios: (1) when both losses and contexts are stochastic, and (2) when losses are adversarial and contexts are stochastic with a known distribution. When losses are adversarial and contexts are stochastic with an unknown distribution, they only achieve the suboptimal $\widetilde{O}(K^{1/3}T^{2/3})$ regret. More recently, Schneider and Zimmert (2023) gave a new algorithm that achieves the nearly optimal $\widetilde{O}(\sqrt{KT})$ regret in expectation under adversarial losses and stochastic contexts with an unknown distribution.
@@
Another interesting application of the cross-learning contextual bandits problem is the sleeping bandits problem (Kleinberg et al., 2010; Neu and Valko, 2014; Kale et al., 2016; Saha et al., 2020). In this problem, a certain set of arms is unavailable in each round. The sleeping bandits problem is motivated by instances like some items might go out of stock in retail stores or on a certain day some websites could be down. When losses are adversarial and availabilities are stochastic, previous work either requires exponential computing time (Kleinberg et al., 2010; Neu and Valko, 2014) or results in suboptimal regret (Kale et al., 2016; Saha et al., 2020). The first computationally efficient algorithm with optimal regret $\widetilde{O}(\sqrt{KT})$ is proposed in Schneider and Zimmert (2023) by modeling the problem as a cross-learning contextual bandit.
@@
2 Problem Statament
We study a contextual $K$-armed bandit problem over $T$ rounds, with contexts belonging to the set $[C]$. At the beginning of the problem, an oblivious adversary selects a sequence of losses $\ell_{t,c}(a)\in[0,1]$ for every round $t\in[T]$, every context $c\in[C]$, and every arm $a\in[K]$. In each round $t$, we begin by sampling a context $c_{t}\sim\nu$ i.i.d. from an unknown distribution $\nu$ over $[C]$, and we reveal this context to the learner. Based on this context, the learner selects an arm $a_{t}\in[K]$ to play. The adversary then reveals the function $\ell_{t,c}(a_{t})$, and the learner suffers loss $\ell_{t,c_{t}}\left(a_{t}\right)$. Notably, the learner observes the loss for every context $c\in[C]$, but only for the arm $a_{t}$ they actually played.
@@ 2 Problem Statament
Schneider and Zimmert (2023) designed an algorithm that achieves an expected regret bound of $\mathbb{E}\left[\operatorname{\mathrm{Reg}}(\pi)\right]\leq\widetilde{O}(\sqrt{KT})$ for any policy $\pi$. We will show that the algorithm in Schneider and Zimmert (2023) actually provides a high-probability regret bound.
3 The Algorithm in Schneider and Zimmert (2023)
In this section, we briefly recap the intuition behind the algorithm proposed in Schneider and Zimmert (2023) and redescribe the algorithm formally to prepare the readers for our new analysis.
3.1 Intuition behind Schneider and Zimmert (2023)
@@
$p_{t,c}=\operatorname*{arg\,min}_{p\in\Delta([K])}\left\langle p,\sum_{s=1}^{t-1}\widehat{\ell}_{s,c}\right\rangle-\frac{1}{\eta}F(p)$
@@
for each context $c$, where $F(p)=\sum_{i=1}^{K}p_{i}\log(p_{i})$ is the unnormalized negative entropy, $\eta$ is a learning rate, and $\widehat{\ell}$ are loss estimates to be defined later. The algorithm then essentially samples the action $a_{t}$ to be played in round $t$ from distribution $p_{t,c_{t}}$.
@@
The key novelty in Schneider and Zimmert (2023) lies in the construction of the loss estimates $\widehat{\ell}$. An intuitive construction is defined as follows:
~$\widetilde{\ell}_{t,c}(a)=\frac{\ell_{t,c}(a)}{\mathbb{E}_{c\sim\nu}[p_{t,c}(a)]}\operatorname{\mathbbm{1}}(a_{t}=a).$
@@
That is, it uses the classic importance-weighted estimator with $\mathbb{E}_{c\sim\nu}[p_{t,c}(a)]$ as the importance [^2] In this paper we call terms like $\frac{1}{\mathbb{E}_{c\sim\nu}[p_{t,c}(a)]}$ as the importance weight and call terms like $\mathbb{E}_{c\sim\nu}[p_{t,c}(a)]$ as the importance.. A straightforward analysis shows that this estimator yields a regret bound of $\widetilde{O}(\sqrt{KT})$. However, the denominator term $\mathbb{E}_{c\sim\nu}[p_{t,c}(a)]$ is uncomputable because we do not know the distribution of contexts $\nu$. One may attempt to circumvent this issue by replacing the expected importance $\mathbb{E}_{c\sim\nu}[p_{t,c}(a)]$ with the empirical importance $\frac{1}{t}\sum_{s=1}^{t}p_{t,c_{s}}(a)$. It is not hard to see that whether we achieve the desired $\widetilde{O}(\sqrt{KT})$ regret depends on how well the empirical importance weight $\frac{1}{\frac{1}{t}\sum_{s=1}^{t}p_{t,c_{s}}(a)}$ concentrates around the expected importance weight $\frac{1}{\mathbb{E}_{c\sim\nu}[p_{t,c}(a)]}$. However, the empirical importance weight $\frac{1}{\frac{1}{t}\sum_{s=1}^{t}p_{t,c_{s}}(a)}$ may not concentrate well around the expected importance weight $\frac{1}{\mathbb{E}_{c\sim\nu}[p_{t,c}(a)]}$. This is because the probability vector $p_{t,c}$ is not independent of the previous contexts $c_{s}$, which makes standard concentration inequalities inapplicable.
@@
To address this difficulty, Schneider and Zimmert (2023) divides the time horizon into epochs of equal length $L$. At the end of each epoch $e$, the algorithm stores the FTRL distribution at the current time $t=eL$ in a new distribution $s_{e}$; that is, it takes $s_{e,c}(a)=p_{t,c}(a)$ for each context $c$ and each arm $a$. The algorithm further decouples the distribution played by the algorithm and the distribution used to estimate the loss vector. For each time $t$ in epoch $e+2$, the algorithm observes the loss $\ell_{t,c}(a)$ for each arm $a$ and context $c$ with probability $f_{e}(a)\triangleq\mathbb{E}_{c\sim\nu}[s_{e,c}(a)/2]$. The algorithm then estimates the expected importance $f_{e}(a)$ using an empirical importance $\widehat{f}_{e}(a)$ constructing solely from contexts in epoch $e+1$. Finally, the algorithm constructs ^$\widehat{\ell}_{t,c}(a)$ as an importance-weighted estimator with $\widehat{f}_{e}(a)$ serving as the importance.
The advantage of their construction is that the empirical importance weight $\frac{1}{\widehat{f}_{e}(a)}$ concentrates around the expected importance weight $\frac{1}{f_{e}(a)}$ now. This concentration ensures that the loss estimates ^$\widehat{\ell}_{t,c}(a)$ are good estimates of the true losses $\ell_{t,c}(a)$. And this concentration is achieved because the algorithm constructs the estimator using only samples from epoch $e+1$, which are independent of the estimand.
3.2 A Formal Description of the Algorithm in Schneider and Zimmert (2023)
@@
In this subsection we describe the algorithm in Schneider and Zimmert (2023) formally for the sake of completeness. Readers familiar with Schneider and Zimmert (2023) can skip this subsection safely.
@@
$p_{t,c}=\operatorname*{arg\,min}_{p\in\Delta([K])}\left\langle p,\sum_{s=1}^{t-1}\widehat{\ell}_{s,c}\right\rangle-\frac{1}{\eta}F(p)$
for each context $c$, where $F(p)=\sum_{i=1}^{K}p_{i}\log(p_{i})$ is the unnormalized negative entropy, $\eta$ is the learning rate, and $\widehat{\ell}$ are loss estimates to be defined later. The algorithm will not sample the action $a_{t}$ played in round t directly from $p_{t}$ but from a distribution $q_{t}$ to be defined later.
To construct loss estimates $\widehat{\ell}$, the algorithm divides the time horizon into epochs of equal length $L$. We let $\mathcal{T}_{e}$ to denote the set of rounds in the $e$-th epoch. At the end of each epoch, the algorithm takes a single snapshot of the underlying FTRL distribution $p_{t}$ for each context and arm. That is, the algorithm takes
$s_{e+2,c}(a)=p_{eL,c}(a)$
and takes
$s_{1,c}(a)=s_{2,c}(a)=\frac{1}{K}.$
The index of the snapshot $s_{e+2}$ for epoch $e$ is $e+2$ because we use the snapshot $s_{e+2}$ in epoch $e+2$.
@@
^$\widehat{\ell}_{t,c}(a)=\frac{2\ell_{t,c}(a)}{\widehat{f}_{e}(a)+\frac{3}{2}\gamma}\operatorname{\mathbbm{1}}\left(A_{t}=a\wedge S_{t}\wedge t\in\mathcal{T}_{e}^{\ell}\right)$
@@
The algorithm is summarized in Algorithm 1. Furthermore, Schneider and Zimmert (2023) showed that the algorithm achieves an expected regret bound of $\widetilde{O}(\sqrt{KT})$.